prettygirl15:

The ratio of roses to daisies is 7 roses to 6 daisies. If there are 21 roses, how many daisies are there?

Nightmoon:

6,12,? figure out ? and you'll find your answer.

ShoaibChaudhry:

That doesnt really help

ShoaibChaudhry:

Ok so listen

ShoaibChaudhry:

7 roses to 6 daisys right

ShoaibChaudhry:

and we have 21 roses deived 21 by 7

ShoaibChaudhry:

which is 3 and then u would time 3 by 6 giving u 18 and thats ur answer :>

jhonyy9:

so just use the rule of 3 simple and write 7 roses to 6 daises 21 roses to x daises cross multiplie and calcule the x value 21*6 = 7*x divide both sides by 7 21*6/7 = x x = ?

jhonyy9:

do you understand the way ?

Her0Hottie:

18

jhonyy9:

@her0hottie wrote:
18
yes of cours